1. Calculate the magnitude and direction of the electric field 0.50m from a +8.23^-9C point charge.

2. An electric field dipole consists of two equal but unlike charges separated by a distance. Two point charges, q1=+2.3×10^-6C and q2=3.4×10^-6, are separated by 8.9×10^-2m, forming an electric dipole. Find the electric field halfway between the dipole.


1
Expert's answer
2022-04-15T16:15:03-0400

1.

"E=\frac{kq}{r^2}=300~\frac NC,"

2.

"E=\frac{4k(q_1+q_2)}{r^2}=26~\frac{MN}C."